OpenStudy (anonymous):

The midpoint of a line segment with end points as (x, 4) and (-2, -2) is (2, 1). What is the value of x? Answer x = -3 x = -2 x = 4 x = 6

OpenStudy (anonymous):

6

OpenStudy (lgbasallote):

\[\large \frac{(x_1 + x_2)}{2} = m_x\] mx means x coordinate of the midpoint

OpenStudy (anonymous):

((x-2)/2,1) x-2/1 =2 x=6

OpenStudy (lgbasallote):

your x1 is x and your x2 = -2 mx = 2 therefore.... \[\frac{x + (-2)}{2} = 2\] \[\frac{x -2}{2} = 2\] cross multiply \[x - 2 = 4\] solve for x

OpenStudy (anonymous):

x-2/2 =2 sorry not divided by one...what he said ^^^

OpenStudy (anonymous):

thanks:)